(1+ix÷1-ix)=a-ib, a2+b2=1 find the value of x .it is complex no topic please answer this question kindly.

Grade:12th pass

1 Answers

Arun
5 years ago
 
Dear Manoj
 
1 + ix /(1- ix) = 1 – x^2 + i * 2x /2 = a – ib
 
hence
a = 1- x^ 2 /2
b = – x
 
now use given relation